# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix A Roadmap of the Proofs
In all four of our main results, the threshold consists of a maximum of two terms, the second of which itself consists of an optimization of parameters $(C,\zeta)$. For both the achievability and converse results, these two terms are obtained using very different methods. Forward-references to the relevant subsequent appendices will be given below.
@@
Achievability results. For the achievability results, we recall the hybrid decoder introduced in (21)–(22) in Section 3.2, which for a given candidate set $s$, does two different checks that correspond to considering $-\ell$ and $-\ell$ errors separately:
@@
  • The second check (corresponding to high $\ell$) is based on thresholding the information density (introduced in (18)), and we study its conditions for success by following the framework of [39] (but with significant gaps to fill in the case of the near-constant weight design).
@@
    The true defective set (say $s=\{1,\dotsc,k\}$ without loss of generality) satisfies (21)–(22);
@@
To see that these conditions are sufficient, first consider some incorrect estimate $\tilde{s}$ with $|s\setminus\tilde{s}|\leq\frac{k}{\log k}$. Such $\tilde{s}$ is ruled out by the $-\ell$ part of our decoder (i.e., (21)), because the correct $s$ is within distance $\frac{k}{\log k}$ and has a higher likelihood due to condition 1 above. On the other hand, for any incorrect estimate $\tilde{s}$ with $|s\setminus\tilde{s}|>\frac{k}{\log k}$, condition 2 above immediately implies that the $-\ell$ condition of the decoder (i.e., (22)) is not satisfied for $\tilde{s}$. Thus, we conclude that the true defective set $s$ satisfies (21)–(22) but none of the incorrect sets $\tilde{s}$ do so, as desired.
@@ A.2 Note on the Scaling of $n$
In our achievability results we seek to show that then error probability approaches zero when $n\geq(1+\eta)n^{*}$, and the in converse results we seek to show that any algorithm has error probability approaching one when $n\leq(1-\eta)n^{*}$, where $n^{*}$ is a threshold scaling as $\Theta(k\log p)$. Recall that throughout the entire paper, we consider the regime $k=\Theta(p^{\theta})$ with $\theta\in(0,1)$, which implies that $\log k=\Theta(\log p)$. This implies that $n^{*}=\Theta(k\log p)=\Theta(k\log k)$ in all of our results.
In view of these observations, we will assume throughout the analysis that $n$ itself also scales as $\Theta(k\log p)=\Theta(k\log k)$. For the achievability part, this is justified by the fact that only $n=\Theta(k\log p)$ or $n=\omega(k\log p)$ are valid choices, and the former clearly provides a stronger result (because of requiring fewer tests). Similar, for the converse part, only $n=\Theta(k\log p)$ or $n=o(k\log p)$ are valid, and a converse for the former immediately implies a converse for the latter anyway (since the decoder could always choose to ignore some of the test results).
@@ A.3 Useful Technical Lemmas
    $\mathbb{P}(X\leq k)\leq\exp\Big{(}-N\cdot D\Big{(}\frac{k}{N}\big{\|}q\Big{)}\Big{)},$ (23)
    which implies
    $\mathbb{P}(X\leq k)\leq\exp\Big{(}-Nq\Big{(}\frac{k}{Nq}\log\frac{k}{Nq}+\frac{k}{Nq}-1\Big{)}\Big{)}.$ (24)
    These bounds also remain true when replacing $\mathbb{P}(X\leq k)$ with $\mathbb{P}(X\geq k)$ for $k\geq Nq$.
@@
    $\mathbb{P}(X=k)\geq\frac{1}{2\sqrt{2k(1-\frac{k}{N})}}\exp\Big{(}-N\cdot D\Big{(}\frac{k}{N}\big{\|}q\Big{)}\Big{)},$ (25)
@@
    $\mathbb{P}(X=k)\geq\frac{1}{\sqrt{2N}}\exp\Big{(}-N\cdot D\Big{(}\frac{k}{N}\big{\|}q\Big{)}\Big{)}.$ (26)
@@
$\frac{\sqrt{\pi}}{2}\cdot\frac{\exp(NH_{2}(k/N))}{\sqrt{2\pi k(1-k/N)}}\leq\binom{N}{k}\leq\frac{\exp(NH_{2}(k/N))}{\sqrt{2\pi k(1-k/N)}}.$ (27)
@@
$\frac{\exp(NH_{2}(k/N))}{\sqrt{2N}}\leq\binom{N}{k}\leq\exp(NH_{2}(k/N)).$ (28)
@@ A.4 Table of Notation
$\rho,\nu,\theta$ Noise parameter, design parameter, sparsity parameter $(k=\Theta(p^{\theta})$)
@@
$d,d^{*}$ Paramter associated with $g$, its optimal value
@@ Appendix B Necessary and Sufficient Conditions for Maximum-Likelihood Decoding
$\widehat{S}=\text{arg\leavevmode\nobreak\ max}_{s^{\prime}\in\mathcal{S}_{k}}\mathcal{L}(s^{\prime}),\text{ where }\mathcal{L}(S):=\mathbb{P}(\mathbf{Y}|S,\mathbf{X}),$ (29)
@@
In our achievability analysis, in view of the two sufficient conditions identified in Appendix A, we will consider a restricted MLE decoder such that the argmax in (29) is restricted to sets within distance $\frac{k}{\log k}$ of the true one. Note that this does not imply that our decoder (introduced in Section 3.2) actually has knowledge of the true defective set; rather, restricted MLE is a hypothetical decoder whose success coincides with one of the sufficient conditions identified in Appendix A. Thus, when the true defective set is $s$, we are interested in the following:
$\widehat{S}^{\prime}=\text{arg\leavevmode\nobreak\ max}_{s^{\prime}\in\mathcal{S}_{k}\,:\,|s\setminus s^{\prime}|\leq\frac{k}{\log k}}\mathcal{L}(s^{\prime}).$ (30)
In fact, it will suffice to also consider this restricted problem in our converse analysis, observing that if $\widehat{S}^{\prime}\neq s$ in (30) then we clearly also have $\widehat{S}\neq s$ in (29).
Under some defective set $S\subset[p]$ with $|S|=k$, we say a test is correct if its result is consistent with the noiseless setting, and for the given $(\mathbf{X},\mathbf{Y})$ we define $N_{\mathbf{X},\mathbf{Y}}(S)$ as the number of correct tests. Observe that
$\mathcal{L}(S)=\rho^{n-N_{\mathbf{X},\mathbf{Y}}(S)}(1-\rho)^{N_{\mathbf{X},\mathbf{Y}}(S)},$ (31)
so due to the fact that $\rho\in\big{(}0,\frac{1}{2}\big{)}$, the MLE decoder finds $\widehat{S}$ with the most correct tests, i.e., $N_{\mathbf{X},\mathbf{Y}}(\widehat{S})\geq N_{\mathbf{X},\mathbf{Y}}(S)$ holds for any $S\in\mathcal{S}_{k}$ (or similarly for $\widehat{S}^{\prime}$ in the restricted version).
@@
In the following lemma, we provide deteriminstic conditions for the failure of the restricted MLE decoder.
Lemma 3.
(Conditions for the Failure of MLE) Let $s\subset[p]$ with cardinality $k$ be the actual defective set, and consider the restricted MLE decoder in (30). Given $\ell\in[1,\frac{k}{\log k}]$, for a feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$ such that $\frac{Cn\nu e^{-\nu}\ell}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}$ are integers, we define
$\mathcal{K}_{\ell,C,\zeta}=\Big{\{}\mathcal{J}\subset s,|\mathcal{J}|=\ell:M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\Big{\}}$ (32)
and let $k_{\ell,C,\zeta}:=|\mathcal{K}_{\ell,C,\zeta}|$. Given $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$, let $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ be the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ that contain some item from $\mathcal{J}^{\prime}$, and $G_{\mathcal{J},\mathcal{J}^{\prime},2}$ be the number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$ that contain some item from $\mathcal{J}^{\prime}$. Then we have the following two statements:
(a) (Sufficient condition for failure) If for some feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$, $\mathcal{K}_{\ell,C,\zeta}$ is non-empty, and there exist some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$ such that
$G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k},$ (33)
then $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$ with $s^{\prime}:=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$, which implies the failure of the restricted MLE decoder (30).
(b) (Necessary condition for failure) If the restricted MLE decoder (30) fails (i.e., it returns some $s^{\prime}\neq s$ with $|s\setminus s^{\prime}|\in[1,\frac{k}{\log k}]$), then for some feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to $\ell:=|s\setminus\widehat{S}|$, $\mathcal{K}_{\ell,C,\zeta}$ is non-empty, and there exist $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$ such that
$G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}.$ (34)
@@
Recall that we let $N_{\mathbf{X},\mathbf{Y}}(S)$ be the number of correct tests under the defective set $s$, which enters the likelihood via (31). We separately prove the statements (a) and (b) by analyzing the number of correct tests.
Proof of (a): For $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$, we note that the length-$k$ $s^{\prime}:=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$ satisfies $|s\setminus s^{\prime}|=\ell$, and $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$ evidently implies the failure of MLE. Thus we only need to prove $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$, and it suffices to show $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})>N_{\mathbf{X},\mathbf{Y}}(s)$. To this end, we compare the number of correct tests under $s$ and $s^{\prime}$ as follows:
@@
  • Next, we add $\mathcal{J}^{\prime}$ to the current defective set $s\setminus\mathcal{J}$ to arrive at $s^{\prime}=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$. Then, the incorrect tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ become correct if they contain some item from $\mathcal{J}^{\prime}$, while the correct tests in $\mathcal{N}_{00}\cup\mathcal{M}_{j0}$ become incorrect if they contain some item from $\mathcal{J}^{\prime}$. This leads to a gain of $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}$ correct tests.
Therefore, by (33) we have
@@
Proof of (b): If restricted MLE fails and returns a set $s^{\prime}$ with $|s\setminus s^{\prime}|=\ell\in[1,\frac{k}{\log k}]$, then we note from the definition (30) that $\mathcal{L}(s^{\prime})\geq\mathcal{L}(s)$, and thus $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})\geq N_{\mathbf{X},\mathbf{Y}}(s)$. Based on $s^{\prime}$, we construct the feasible $(C,\zeta)$ for which the conclusion of (b) holds. Specifically, we let $\mathcal{J}=s\setminus s^{\prime}$ with $|\mathcal{J}|=\ell$ and recall the definitions of $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1}$ and their cardinalities $M_{\mathcal{J}},M_{\mathcal{J}0},M_{\mathcal{J}1}$; then, we set $(C,\zeta)=(\frac{M_{\mathcal{J}}}{k^{-1}n\nu e^{-\nu}\ell},\frac{M_{\mathcal{J}0}}{M_{\mathcal{J}}})$ if $M_{\mathcal{J}}>0$, or $(C,\zeta)=(0,0)$ if $M_{\mathcal{J}}=0$. By construction $(C,\zeta)\in[0,\infty)\times[0,1]$ is feasible, and by (32) it is easy to see that $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$. Moreover, we let $\mathcal{J}^{\prime}=s^{\prime}\setminus s\subset[p]\setminus s$ satisfying $|\mathcal{J}^{\prime}|=\ell$, and note that by the definition of $(\mathcal{J},\mathcal{J}^{\prime})$ we have $s^{\prime}=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$. It remains to show that (34) holds for the $(\mathcal{J},\mathcal{J}^{\prime})$ we are considering. To this end, we compare the number of correct tests under $s$ and $s^{\prime}$ exactly the same as the above two dot points in the proof of (a): Starting with the true defective set $s$, removing $\mathcal{J}$ from the defective set leads to a loss of $(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}$ correct tests, while adding $\mathcal{J}^{\prime}$ to the current $s\setminus\mathcal{J}$ leads to a gain of $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}$ correct tests. Overall, we obtain $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})-N_{\mathbf{X},\mathbf{Y}}(s)=G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}-(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}$, and thus $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})\geq N_{\mathbf{X},\mathbf{Y}}(s)$ yields (34) as claimed. ∎
@@
In the achievablity analysis, we seek to establish the threshold for $n$ above which the MLE fails with $o(1)$ probability. Since 3(b) provides a necessary condition for the failure of MLE, we can instead bound the probability of the events stated in Lemma 3(b).

### 23 Oct 2024 v2

@@ Appendix A Roadmap of the Proofs
In all four of our main results, the threshold consists of a maximum of two terms, the second of which itself consists of an optimization of the parameters $(C,\zeta)$. In each case, these two terms will arise either from two separate analyses, or from two different error events within the same analysis. Forward-references to the relevant subsequent appendices will be given in Section A.1 below.
@@
Achievability results. For the achievability results, we recall the hybrid decoder introduced in (3.4)–(3.5) in Section 3.2, which for a given candidate set $s$, does two different checks that correspond to considering $-\ell$ and $-\ell$ errors separately:
@@
  • The second check (corresponding to high $\ell$) is based on thresholding the information density (as defined in (3.1)), and we study its conditions for success by following the framework of [39] (but with significant gaps to fill in the case of the near-constant weight design).
@@
    The true defective set (say $s=\{1,\dotsc,k\}$ without loss of generality) satisfies (3.4)–(3.5);
@@
To see that these conditions are sufficient, first consider some incorrect estimate $\tilde{s}$ with $|s\setminus\tilde{s}|\leq\frac{k}{\log k}$. Such $\tilde{s}$ is ruled out by the $-\ell$ part of our decoder (i.e., (3.4)), because the correct $s$ is within distance $\frac{k}{\log k}$ and has a higher likelihood due to condition 1 above. On the other hand, for any incorrect estimate $\tilde{s}$ with $|s\setminus\tilde{s}|>\frac{k}{\log k}$, condition 2 above immediately implies that the $-\ell$ condition of the decoder (i.e., (3.5)) is not satisfied for $\tilde{s}$. Thus, we conclude that the true defective set $s$ satisfies (3.4)–(3.5) but none of the incorrect sets $\tilde{s}$ do so, as desired.
@@ A.2 Note on the Scaling of $n$
Both our achievability and converse results are stated in terms of a threshold $n^{*}$ scaling as $\Theta(k\log p)$. Recall that throughout the entire paper, we consider the regime $k=\Theta(p^{\theta})$ with $\theta\in(0,1)$, which implies that $\log k=\Theta(\log p)$. This implies that $n^{*}=\Theta(k\log p)=\Theta(k\log k)$ in all of our results.
In view of this observation, we will assume throughout the analysis that the number of tests $n$ itself also scales as $\Theta(k\log p)=\Theta(k\log k)$. For the achievability part, this is justified by the fact that the theorems state that we can achieve $P_{\mathrm{e}}\to 0$ while satisfying $n\leq(1+\eta)n^{*}$, and we are free to let this inequality hold with equality or “near-equality” (say $n\in[n^{*},(1+\eta)n^{*}]$). For the converse part, the theorems state that attaining $P_{\mathrm{e}}\not\to 1$ requires $n\geq(1-\eta)n^{*}$, and we will prove the contrapositive statement that $P_{\mathrm{e}}\to 1$ whenever $n<(1-\eta)n^{*}$. The assumption $n=\Theta(k\log p)$ is then justified by noting that if $n=o(k\log p)$, then existing results already show that $P_{\mathrm{e}}\to 1$ even in the case of noiseless tests [16].
@@ A.3 Useful Technical Lemmas
    $\mathbb{P}(X\leq k)\leq\exp\Big{(}-N\cdot D\Big{(}\frac{k}{N}\big{\|}q\Big{)}\Big{)},$ (A.1)
    which remains true when replacing $\mathbb{P}(X\leq k)$ with $\mathbb{P}(X\geq k)$ for $k\geq Nq$. This implies
    $\mathbb{P}(X\leq k)\leq\exp\Big{(}-Nq\Big{(}\frac{k}{Nq}\log\frac{k}{Nq}+1-\frac{k}{Nq}\Big{)}\Big{)}$ (A.2)
    for $k\leq Nq$, which remains true when replacing $\mathbbm{P}(X\leq k)$ with $\mathbb{P}(X\geq k)$ for $k\geq Nq$.
@@
    $\mathbb{P}(X=k)\geq\frac{1}{2\sqrt{2k(1-\frac{k}{N})}}\exp\Big{(}-N\cdot D\Big{(}\frac{k}{N}\big{\|}q\Big{)}\Big{)},$ (A.3)
@@
    $\mathbb{P}(X=k)\geq\frac{1}{\sqrt{2N}}\exp\Big{(}-N\cdot D\Big{(}\frac{k}{N}\big{\|}q\Big{)}\Big{)}.$ (A.4)
@@
$\frac{\sqrt{\pi}}{2}\cdot\frac{\exp(NH_{2}(k/N))}{\sqrt{2\pi k(1-k/N)}}\leq\binom{N}{k}\leq\frac{\exp(NH_{2}(k/N))}{\sqrt{2\pi k(1-k/N)}}.$ (A.5)
@@
$\frac{\exp(NH_{2}(k/N))}{\sqrt{2N}}\leq\binom{N}{k}\leq\exp(NH_{2}(k/N)).$ (A.6)
@@ A.4 Table of Notation
$\rho,\nu,\theta$ Noise level, design parameter, sparsity parameter $(k=\Theta(p^{\theta})$)
@@
$d,d^{*}$ Parameter associated with number of “relevant” positive tests a non-defective is in
@@ Appendix B Necessary and Sufficient Conditions for Maximum-Likelihood Decoding
$\widehat{S}=\text{arg\leavevmode\nobreak\ max}_{s^{\prime}\in\mathcal{S}_{k}}\mathcal{L}(s^{\prime}),\text{ where }\mathcal{L}(S):=\mathbb{P}(\mathbf{Y}|S,\mathbf{X}),$ (B.1)
@@
In our achievability analysis, in view of the two sufficient conditions identified in Appendix A, we will consider a restricted MLE decoder such that the argmax in (B.1) is restricted to sets within distance $\frac{k}{\log k}$ of the true one. Note that this does not imply that our decoder (introduced in Section 3.2) actually has knowledge of the true defective set; rather, restricted MLE is a hypothetical decoder whose success coincides with one of the sufficient conditions identified in Appendix A. Thus, when the true defective set is $s$, we are interested in the following:
$\widehat{S}^{\prime}=\text{arg\leavevmode\nobreak\ max}_{s^{\prime}\in\mathcal{S}_{k}\,:\,|s\setminus s^{\prime}|\leq\frac{k}{\log k}}\mathcal{L}(s^{\prime}).$ (B.2)
In fact, it suffices to also consider this restricted problem in our converse analysis, observing that if $\widehat{S}^{\prime}\neq s$ in (B.2) then we clearly also have $\widehat{S}\neq s$ in (B.1).
For a given set $s^{\prime}\subset[p]$ with $|s^{\prime}|=k$, we say that a test is correct (with respect to $s^{\prime}$) if its result matches what the noiseless result would be under defective set $s^{\prime}$. Given the test matrix and outcomes, $(\mathbf{X},\mathbf{Y})$, we define $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})$ as the number of correct tests. Observe that
$\mathcal{L}(s^{\prime})=\rho^{n-N_{\mathbf{X},\mathbf{Y}}(s^{\prime})}(1-\rho)^{N_{\mathbf{X},\mathbf{Y}}(s^{\prime})},$ (B.3)
so due to the fact that $\rho\in\big{(}0,\frac{1}{2}\big{)}$, the MLE decoder finds $\widehat{S}$ with the most correct tests, i.e., $N_{\mathbf{X},\mathbf{Y}}(\widehat{S})\geq N_{\mathbf{X},\mathbf{Y}}(s^{\prime})$ holds for any $s^{\prime}\in\mathcal{S}_{k}$ (or similarly for $\widehat{S}^{\prime}$ in the restricted version).
@@
In the following lemma, we provide deterministic conditions for the failure of the restricted MLE decoder.
Lemma B.1.
(Conditions for the Failure of MLE) Let $s\subset[p]$ with cardinality $k$ be the actual defective set, and consider the restricted MLE decoder in (B.2). Given $\ell\in[1,\frac{k}{\log k}]$, for a feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$ such that $\frac{Cn\nu e^{-\nu}\ell}{k}$ and $\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}$ are integers, we define
$\mathcal{K}_{\ell,C,\zeta}=\Big{\{}\mathcal{J}\subset s\,:\,|\mathcal{J}|=\ell,M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\Big{\}}$ (B.4)
and let $k_{\ell,C,\zeta}:=|\mathcal{K}_{\ell,C,\zeta}|$. Given $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$, let $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ be the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ that contain some item from $\mathcal{J}^{\prime}$, and let $G_{\mathcal{J},\mathcal{J}^{\prime},2}$ be the number of tests in $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$ that contain some item from $\mathcal{J}^{\prime}$. Then we have the following two statements:
(a) (Sufficient condition for failure) If for some feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to some $\ell\in[1,\frac{k}{\log k}]$, the set $\mathcal{K}_{\ell,C,\zeta}$ is non-empty, and there exist some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$ such that
$G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}>(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k},$ (B.5)
then $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$ with $s^{\prime}:=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$, which implies the failure of the restricted MLE decoder (B.2).
(b) (Necessary condition for failure) If the restricted MLE decoder (B.2) fails (i.e., it returns $\widehat{S}=s^{\prime}$ for some $s^{\prime}\neq s$ with $|s\setminus s^{\prime}|\in[1,\frac{k}{\log k}]$), then for some feasible pair $(C,\zeta)\in[0,\infty)\times[0,1]$ with respect to $\ell:=|s\setminus\widehat{S}|$, the set $\mathcal{K}_{\ell,C,\zeta}$ is non-empty, and there exist some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$ such that
$G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}.$ (B.6)
@@
Recall that we let $N_{\mathbf{X},\mathbf{Y}}(s)$ be the number of correct tests under the defective set $s$, which enters the likelihood via (B.3). We separately prove the statements (a) and (b) by analyzing the number of correct tests.
Proof of (a): For $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$, we note that the size-$k$ set $s^{\prime}:=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$ satisfies $|s\setminus s^{\prime}|=\ell$, and $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$ evidently implies the failure of MLE. Thus we only need to prove $\mathcal{L}(s^{\prime})>\mathcal{L}(s)$, and it suffices to show $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})>N_{\mathbf{X},\mathbf{Y}}(s)$. To this end, we compare the number of correct tests under $s$ and $s^{\prime}$ as follows:
@@
  • Next, we add $\mathcal{J}^{\prime}$ to $s\setminus\mathcal{J}$ to arrive at $s^{\prime}=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$. Then, the incorrect tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ become correct if they contain some item from $\mathcal{J}^{\prime}$, while the correct tests in $\mathcal{N}_{00}\cup\mathcal{M}_{j0}$ become incorrect if they contain some item from $\mathcal{J}^{\prime}$. This leads to a gain of $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}$ correct tests.
Therefore, by (B.5) we have
@@
Proof of (b): If restricted MLE fails and returns a set $s^{\prime}$ with $|s\setminus s^{\prime}|=\ell\in[1,\frac{k}{\log k}]$, then we note from the definition (B.2) that $\mathcal{L}(s^{\prime})\geq\mathcal{L}(s)$, and thus $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})\geq N_{\mathbf{X},\mathbf{Y}}(s)$. Based on $s^{\prime}$, we construct the feasible $(C,\zeta)$ for which the conclusion of (b) holds. Specifically, we let $\mathcal{J}=s\setminus s^{\prime}$ with $|\mathcal{J}|=\ell$ and recall the definitions of $\mathcal{M}_{\mathcal{J}},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1}$ and their cardinalities $M_{\mathcal{J}},M_{\mathcal{J}0},M_{\mathcal{J}1}$; then, we set $(C,\zeta)=(\frac{M_{\mathcal{J}}}{k^{-1}n\nu e^{-\nu}\ell},\frac{M_{\mathcal{J}0}}{M_{\mathcal{J}}})$ if $M_{\mathcal{J}}>0$, or $(C,\zeta)=(0,0)$ if $M_{\mathcal{J}}=0$. By construction $(C,\zeta)\in[0,\infty)\times[0,1]$ is feasible, and by (B.4) it is easy to see that $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$. Moreover, we let $\mathcal{J}^{\prime}=s^{\prime}\setminus s\subset[p]\setminus s$ satisfying $|\mathcal{J}^{\prime}|=\ell$, and note that by the definition of $(\mathcal{J},\mathcal{J}^{\prime})$ we have $s^{\prime}=(s\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$. It remains to show that (B.6) holds for the $(\mathcal{J},\mathcal{J}^{\prime})$ we are considering. To this end, we compare the number of correct tests under $s$ and $s^{\prime}$ exactly the same as the above two dot points in the proof of (a): Starting with the true defective set $s$, removing $\mathcal{J}$ from the defective set leads to a loss of $(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}$ correct tests, while adding $\mathcal{J}^{\prime}$ to the current $s\setminus\mathcal{J}$ leads to a gain of $G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}$ correct tests. Overall, we obtain $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})-N_{\mathbf{X},\mathbf{Y}}(s)=G_{\mathcal{J},\mathcal{J}^{\prime},1}-G_{\mathcal{J},\mathcal{J}^{\prime},2}-(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}$, and thus $N_{\mathbf{X},\mathbf{Y}}(s^{\prime})\geq N_{\mathbf{X},\mathbf{Y}}(s)$ yields (B.6) as claimed. ∎
@@
In the achievability analysis, we seek to establish the threshold for $n$ above which the MLE fails with $o(1)$ probability. Since B.1(b) provides a necessary condition for the failure of MLE, we can instead bound the probability of the events stated in Lemma B.1(b).